A wire is 10 meters long is divided into 3 parts, one of which is 2 meters longer than the first and the third is 2 times longer than the second. Find the length of each part?
Answer by edjones(7311)

Let the length of the 1st part be x.
x, x+2 and 2(x+2) are the length of the 3 parts.
x+x+2+2(x+2)=10
2x+2+2x+4=10
4x+6=10
4x=4
x=1 m
So the lengths are 1, 3, 6 m
.
Ed